# Information Elicitation in Agency Games

### v2

@@ Abstract
Rapid progress in scalable, commoditized tools for data collection and data processing has made it possible for firms and policymakers to employ ever more complex metrics as guides for decision-making. These developments have highlighted a prevailing challenge—deciding which metrics to compute. In particular, a firm’s ability to compute a wider range of existing metrics does not address the problem of unknown unknowns, which reflects informational limitations on the part of the firm. To guide the choice of metrics in the face of this informational problem, we turn to the evaluated agents themselves, who may have more information than a principal about how to measure outcomes effectively. We model this interaction as a simple agency game, where we ask: When does an agent have an incentive to reveal the observability of a cost-correlated variable to the principal? There are two effects: better information reduces the agent’s information rents but also makes some projects go forward that otherwise would fail. We show that the agent prefers to reveal information that exposes a strong enough differentiation between high and low costs. Expanding the agent’s action space to include the ability to garble their information, we show that the agent often prefers to garble over full revelation. Still, giving the agent the ability to garble can lead to higher total welfare. Our model has analogies with price discrimination, and we leverage some of these synergies to analyze total welfare.
@@ 1 Introduction
The rise of algorithmic and data-driven decision-making has come with an accompanying rise in reliance on numerical measurements of performance. These numerical metrics drive actions at scales ranging from that of individual workers such as teachers [Koretz, 2017], to entire institutions such as hospitals [Muller, 2019], and can have far-reaching impact on social welfare. The problem of designing and choosing evaluation metrics thus continues to be both highly consequential and challenging for all types of organizations, from government agencies to companies.
For example, consider the evaluation-metric design problem faced by a company with a research division. To monitor the health of the organization and guide compensation, metrics are collected to evaluate the research division’s productivity. To choose these metrics, company leadership may start with metrics obtained from ad hoc brainstorming or prior experience, taking the form of simple measures such as the number of academic publications, patents, conference presentations, and citations to all of these, along with collaborations with product teams. Such measures may not, however, fully capture the performance of the research division—in fact, they will almost inevitably omit some important factors. The challenge in designing a monitoring system is thus not only in developing methods for measuring performance and optimizing a given set of metrics, but more broadly how to bring new metrics to the fore that were previously unknown and are highly relevant.
This frames our central question of how to improve a firm’s specification of metrics. The perspective underlying this work is that while company leadership may operate under incomplete information, this does not mean that better information does not exist elsewhere in the firm. In fact, often, agents being evaluated may have better information not only about their effort, but also about how to measure outcomes in a more effective, less gameable way. In the above example, researchers might know of other metrics capturing the quality of their work that company leadership had overlooked, like the comparative quality of publication venues. Thus, we model and analyze situations under which an agent being evaluated might be willing to share information with the principal to improve the metrics that are used for evaluation. A key strategic dependence in this work is the relationship between the informant (the agent) and the metrics—we consider an agent’s incentives to reveal information about metrics that are directly used to evaluate them.
@@
The incomplete nature of metrics has been discussed in the seminal work of Holmström and Milgrom [1991] on contracts with multidimensional tasks, where a principal may observe only a subset of dimensions that are relevant to their value or the agent’s cost. When the principal is aware of which dimensions are missing, Holmström and Milgrom [1991] show how to optimally reward the observed dimensions given properties of the agent’s cost structures. The distinguishing feature of our setting is that we consider the unobserved dimensions to be unknown unknowns. We focus on an information transfer mechanism where the agent has the power to possibly reveal these hidden dimensions to the principal.
Our question also fits into a broad class of problems on information design and persuasion games that studies the welfare effects of different information structures [Kamenica and Gentzkow, 2011, Milgrom and Roberts, 1986]. Our model treats an agent as a sender and principal as a receiver, and can be seen as applying specific restrictions to the class of information-revelation strategies that the agent can choose from in order to capture properties of the design of metrics in agency problems faced by firms. Specifically, we consider a setting where the agent’s information revelation strategy cannot depend on the value of the realized signal. Thus, we seek to model an agent’s choice of whether or not to reveal the observability of a metric, rather than a realized signal value.
@@
To yield initial tractable insights to this metric discovery problem, we examine an agent’s incentives for information sharing through the lens of an agency game with information transfer. We build on a classical agency game where a principal contracts an agent to complete a task, and the principal only has partial information about the agent’s costs when setting a contract. To capture the agent’s additional information and opportunity to improve the metrics by which they are evaluated, our model supposes that the agent is privately aware of additional variables that correlate with their cost of task completion, and further has the opportunity to reveal these additional variables to the principal prior to the design of the contract. We analyze when the agent would prefer for the contract to depend on these cost-correlated variables compared to an agnostic contract. Importantly, the agent must decide whether to reveal their cost-correlated variables to the principal prior to realizing the values of those variables or their true costs.
@@
While the principal is always better off when having more information about the agent’s costs, the incentives for the agent are more nuanced—revealing information reduces the amount of information rent the agent can extract when their costs turn out to be low. However, better information also means more high-cost jobs will go forward that the agent otherwise would not have accepted, as the optimal contract would adequately reward high-cost tasks.
We first consider whether the agent prefers to reveal or conceal the observability of a cost-correlated environmental variable. We show that the agent prefers to reveal the variable if conditioning on this cost-correlated variable reveals a strong enough differentiation between high and low costs of task completion. Next, we expand the agent’s action space to include the ability to garble or add noise to their information before it passes to the principal. For example, in modern online platforms, avenues for introducing this type of noise include third party clients or policies requiring differential privacy [Dwork et al., 2006]. We show that under a fairly wide set of conditions, the agent may prefer to reveal a garbled signal over both fully concealing and fully revealing the original variable. This suggests that a noisy information transfer mechanism can yield superior equilibria over simpler or more restrictive alternatives.
1.1 Contributions
@@
    We introduce a model for the process of discovery of unknown metrics, in the form of an agency game with information transfer.
@@
    We present sufficient conditions under which an agent would prefer revealing a metric to a principal over keeping the metric concealed; and vice versa.
@@
    We further relax the agent’s action space to include the ability to reveal garbled information, where the choice of the amount of garbling interpolates between concealment and revelation. In this setting, we give sufficient conditions for the agent to prefer garbling over full revelation.
@@
    We analyze the consequences of information revelation on principal utility and total welfare, leveraging connections between our model and price discrimination.
1.2 Related Work
Our model builds on literature from contract design and agency games. It also fits into a large literature on information design, and can be seen as a specific structure of information design problem. Our model also overlaps with price discrimination, and we bring techniques from these literatures to a new motivation of discovering metrics.
Agency games and contract design
We build from the well-established contract design problem of Laffont and Tirole [1986], which concerns a principal’s design of a contract when an agent’s effort and cost type are privately held by the agent. Key to this setting are asymmetries in values and information between the principal and agent, and the literature explores issues of moral hazard and adverse selection that arise from these asymmetries [Laffont and Martimort, 2009, Gibbons et al., 2013, Milgrom and Roberts, 1992].
@@ 1.2 Related Work
The use of stakeholder-supplied information for decision-making was seminally introduced by Milgrom and Roberts [1986] in the form of persuasion games. Milgrom and Roberts [1986] give an example of understanding a buyer’s purchasing strategy in the face of a seller who can send a quality signal about their product. While this broad motivation of information transfer from interested parties is very close to our motivation of learning business metrics from evaluated agents, the game that we present differs in two main ways from the particular setting in Milgrom and Roberts [1986].
First, we restrict the agent’s information revelation strategy for a given variable to not be able to depend on the realized value of the variable. In other words, the agent must commit to fully revealing a variable in advance of realizing the variable’s value. In this sense, our setting can be viewed as a constrained persuasion game. The purpose of this restriction is to bring our setting closer to a problem of variable discovery, where the fundamental problem is a principal’s lack of awareness of a variable’s observability, rather than signaling [Spence, 1978], where the uncertainty is about the variable’s realized value.
The second main distinction is in the direction of information transfer. In our setting, the “decision-maker” is the agent, who decides whether or not to complete a task. We consider the agent’s incentive to reveal information about their decision-making parameters (their cost type) to a principal who designs a payment contract. This second distinction is more technical and less important for the fundamental motivation of revealing observability instead of value.
@@
More broadly, Bayesian persuasion and information design provides a general framework for analyzing the effects of the distribution of information on the outcomes of a game [Kamenica, 2019, Bergemann and Morris, 2019]. Many works have cast this general form to analyze applications from grading systems in schools to courtroom evidence policies [Kamenica and Gentzkow, 2011, Boleslavsky and Cotton, 2015, Ostrovsky and Schwarz, 2010]. Our model can be seen as a specific instantiation of an information design problem where the sender is the agent, the receiver is the principal, and the contracting relationship determines the principal’s and agent’s action spaces and equilibria. We impose a constraint over the sender’s information transfer policy in order to capture the incentives for an agent to reveal observability of a variable to the principal, whereas prior work has focused on revealing the value of a signal (similarly to Milgrom and Roberts [1986]). Also built into this constraint is that the agent cannot lie about their signal, as our model is motivated by understanding settings where a principal has powerful data collection and verification capabilities.
Price discrimination
@@
Our analysis framework of an agency game with information transfer has analogies with classical price discrimination, and thereby opens new avenues for applying the well-developed tools from price discrimination to understanding the metric design problem. It also uncovers new challenges that extend existing price discrimination perspectives. This connection has also been discussed by Bergemann et al. [2015], who define a mapping from third degree price discrimination onto the class of agency problems and establish the existence of market segmentations that achieve all possible trade-offs between consumer and producer surplus within some basic constraints. Our work complements this analysis—instead of analyzing all possible segmentations, we consider the agent’s information-revelation incentives as a function of the properties of a specific segmentation induced by some cost-correlated variable which is initially only available to the agent.
@@
Our garbling setting also notably differs from the models of price discrimination with transportation costs and arbitrage [Wright, 1993] or restricted price discrimination [Aguirre et al., 2010]. While both garbling and restricted price discrimination mechanisms effectively interpolate between full discrimination and no discrimination, we prove a substantive difference between these mechanisms by showing that while the agent might prefer some intermediate amount of garbling over both full revelation or hiding, the agent will always prefer either full discrimination or no discrimination over any intermediate restriction over the amount of allowed price discrimination. We give a more detailed comparison of our garbling model to price discrimination in Section 4.7.
Sunspots and correlated equilibria
@@
Our model is also connected to the so-called sunspots literature [Woodford, 1990, Howitt and McAfee, 1992]. In this literature, there are multiple steady-state equilibria and an observable random variable produces a correlated equilibrium, with agents conditioning their behavior on this otherwise extraneous variable not because it matters to payoffs, but because it predicts others’ behaviors. The literature is known as sunspots because, during the 19th century, some people believed sunspot activity predicted agricultural yields [Jevons, 1884], and while it did not, it could predict the behavior of commodity traders who believed it did. This literature is the polar opposite of the problem we study: sunspots are a known, extraneous variable believed to be relevant, while we study an unknown (to the principal) relevant variable.
Preferences for privacy
@@
Our model also interacts with literatures on privacy and information-hiding, by highlighting a simple situation where agents experience conflicting incentives both for and against sharing information. The result can be that agents may prefer partial sharing, which echoes the subtleties that arise elsewhere when privacy and behavior interact [Cummings et al., 2016].
2 Model Setup: Agency Game with Information Transfer
@@
Specifically, suppose a principal contracts the agent to complete a task, where an agent may exert binary effort. Suppose the principal receives value $b$ if the agent exerts effort and completes the task, and zero otherwise (we assume that the task is completed deterministically if the agent exerts effort). Suppose the agent incurs cost $C\in\mathbb{R}_{+}$ for exerting effort. The exact cost is unobserved to the principal, but the principal is aware of a prior distribution over agent’s cost type, denoted by the random variable $C$. The agent observes both the proposed transfer and their realized cost type before deciding whether or not to exert effort. This maps onto the well-established agency game setup where the principal must design a contract when the agent’s cost and effort are private [Laffont and Tirole, 1986].
To model a setting where the agent might possess additional information, suppose the agent is aware of an environmental variable $X\in\mathcal{X}$ which is correlated with cost $C$. The question of knowledge of $X$ becomes significant when $X$ is correlated with $C$.
In introducing this correlation into the agency game, we assume that $X$ is an unknown unknown to the principal. Prior to the principal’s design of the contract, the agent has a choice of whether or not to inform the principal of the environmental variable $X$, which entails revealing both the observability of $X$ at the time of the design of the contract, and the realized value of $X$ at the execution of the contract (which the principal can verify). If the agent chooses not to inform the principal of the observability of $X$, then the rest of the game proceeds as a standard agency game with private cost: the principal designs a contract based on their knowledge of the prior distribution $C$. If the agent chooses to inform the principal of $X$, then the principal can offer a contract that conditions on the realized value of $X$.
The timing of the full agency game with the possibility of agent information transfer is summarized in Figure 1. The text in black matches the standard agency game [Laffont and Tirole, 1986], and the text in gray represents additional elements introduced by our model.
@@
In our simplified environment involving binary effort, the principal’s contract design problem when $X$ is concealed reduces to choosing a single transfer amount $p$ where the agent is paid $p$ if the task is completed, and zero otherwise. If $X$ is revealed, then the principal offers a contract with distinct transfers $\rho(x)$ for different realized values of $x\in\mathcal{X}$. At the time of the execution of the contract, the agent receives transfer $\rho(x)$ if the task is completed and $X=x$; and zero otherwise. We assume that the principal still receives the same value $b$ if the task is completed, regardless of $X$.
The key question in this work concerns the agent’s decision of whether or not to inform the principal of the existence of the environmental variable $X$ at time $t=1$. In Section 3, we consider this as a binary decision of whether or not to reveal $X$; we will later relax this in Section 4 to expand the agent’s action space to reveal a garbled version of $X$, thus interpolating between the concealed and revealed settings.
2.1 Notation
Let $F(c)={\mathbb{P}}(C\leq c)$ denote the cumulative distribution function (CDF) of the prior cost distribution $C$. We also define $F_{x}(c)={\mathbb{P}}(C\leq c|X=x)$ as the CDF of the conditional distribution of $C$ given $X$. Throughout, we will assume that the cost $C$ is a continuous random variable with density $f(c)=F^{\prime}(c)$ and with conditional density $f_{x}(c)=F_{x}^{\prime}(c)$.
We will use $\Pi$ to denote the principal’s utility, $V$ to denote the agent’s utility, and $W$ to denote total welfare, defined as the sum of the principal’s and agent’s utilities.
Let $\operatorname{\mathbbm{1}}(\cdot)$ denote an indicator function with $\operatorname{\mathbbm{1}}(x\in S)=\begin{cases}1&\text{if }x\in S\\ 0&\text{otherwise}.\end{cases}$
All proofs are given in the Appendix.

### v3

# Relying on the Metrics of Evaluated Agents

@@ Abstract
Online platforms and regulators face a continuing problem of designing effective evaluation metrics. While tools for collecting and processing data continue to progress, this has not addressed the problem of unknown unknowns, or fundamental informational limitations on part of the evaluator. To guide the choice of metrics in the face of this informational problem, we turn to the evaluated agents themselves, who may have more information about how to measure their own outcomes. We model this interaction as an agency game, where we ask: When does an agent have an incentive to reveal the observability of a metric to their evaluator? We show that an agent will prefer to reveal metrics that differentiate the most difficult tasks from the rest, and conceal metrics that differentiate the easiest. We further show that the agent can prefer to reveal a metric garbled with noise over both fully concealing and fully revealing. This indicates an economic value to privacy that yields Pareto improvement for both the agent and evaluator. We demonstrate these findings on data from online rideshare platforms.
@@ 1 Introduction
“Only the wearer knows where the shoe pinches.” –Proverb
The design of effective evaluation metrics continues to be of central concern to online platforms and their surrounding economies, including regulators, engineers, and users. Indeed, the problem has only grown in importance in recent years, as advances in the ability of online platforms to collect and analyze data have led to a reliance on data-driven evaluation metrics to steer decision-making. The design problem becomes an adaptive and incentive-theoretic one.
For instance, in the rapidly evolving AI marketplace, government regulators are increasingly concerned with the evaluation of AI systems, focusing attention on the evaluation of safety, performance, and liability, in the context of evolving law and policy [House, 2023]. Private companies have demonstrated investment in improving evaluation as well, from developing leaderboards [Fourrier et al., 2024, Scale AI, 2024] to funding evaluation efforts from third-party researchers [Anthropic, 2024]. Academic researchers are at the forefront of efforts to develop new evaluation metrics that target an ever-increasing spectrum of capabilities [Ghazal et al., 2013, Liang et al., 2022].
Within the firms developing these technologies, evaluation has also long been a core part of operations, where internal evaluation metrics drive company decisions in all areas, from engineering to advertising to pricing. Online platforms (like recommender systems, digital marketplaces, and gig economy platforms) draw from rich data sources to compute their internal evaluation metrics. Still, significant challenges remain in the design and usage of these metrics, from handling the limitations and biases of current “proxy” metrics [Athey et al., 2019, Tripuraneni et al., 2024], to identifying which aspects of the data are most relevant to a given decision [Muller, 2019, Bloomfield, 2016].
We consider the development of evaluation metrics, where we use the word “metric” to loosely refer to a variable that an evaluator uses to infer the value or difficulty of a task under consideration. For example, a score on an evaluation benchmark is a metric that indicates the value of a LLM trained by a firm. The delivery time is a metric that indicates both value and difficulty of a food delivery completed by a driver. The h-index is a metric that some employers may use to infer the productivity of a researcher. Metrics almost never tell the whole story, and vary in informativeness, but are nonetheless deeply embedded in online platforms and economies, driving the evolution of technology and shaping its societal impact.
For both firms and regulators, the modern prevailing challenge with the design of metrics is not necessarily with the ability to collect large volumes of data, but of discerning which aspects of the data are relevant to the task at hand. This knowledge in turn informs allocation of resources for data-collection efforts, and downstream decisions that affect users like pricing and auditing. We consider evaluators who are powerful, but with a limited worldview: they can verify and collect any data for metrics that they are aware of, but their worldview is limited in that they are not aware of all possible metrics that could be important.
A core perspective underlying this work is that while evaluators might be limited by their worldviews, this does not mean that better information does not exist in the wider socio-technical system. In fact, those who know the most about a task are often none other than the agents who perform it themselves. These evaluated agents might both know and be willing to share better evaluation metrics. Thus, in this work, we study the incentives for information transfer from an evaluated agent to the evaluator.
More generally, there is a long history of eliciting information and feedback from the wider socio-technical system for the purposes of evaluation and design. For example, the subfield of participatory design in human-computer interaction focuses on methods for designing technical systems with user input [Schuler and Namioka, 1993]. For AI evaluation, some of the biggest companies currently developing the AI systems are also involved in the evaluation process, even cited as direct collaborators in a recent White House briefing [House, 2023]. In online content and labor platforms, there is often a rich reliance on feedback channels like surveys and focus groups of creators and workers on the platform.
We present a theoretical model of incentives for an agent to share metrics across such channels. We characterize the incentives that influence an evaluated agent to share metrics with an evaluator, with a particular focus on the unique capabilities and limitations of modern online platforms. Specifically, we consider a setting in which there are significant asymmetries in data collection, data verification, and price-setting power between the evaluator and the evaluated agents. Yet, a key information asymmetry remains in the other direction, where the evaluated agents hold unique information about what features are most correlated with their success. A core strategic dependence in our model is the relationship between the agent and the metrics—we consider the incentives an agent may experience with respect to the revelation of metrics that impact their own evaluation by an outside entity.
@@
The incomplete nature of metrics has been discussed in the seminal work of Holmström and Milgrom [1991] on contracts with multidimensional tasks, where a principal may observe only a subset of dimensions that are relevant to their value or the agent’s cost. When the principal is aware of which dimensions are missing, Holmström and Milgrom [1991] show how to optimally reward the observed dimensions given properties of the agent’s cost structures. The distinguishing feature of our setting is that we consider the unobserved dimensions to be unknown unknowns, yielding a model that is in a sense inherently non-Bayesian. We focus on an information-transfer mechanism where the agent has the power to possibly reveal these hidden dimensions to the principal.
@@
Specifically, we examine an agent’s incentives for information sharing through the lens of an agency game with information transfer. We build on a classical agency game where a principal contracts an agent to complete a task, and the principal only has partial information about the agent’s costs [^1] In addition to representing the difficulty of the task, the agent’s “cost” could also be interpreted as the external market value for the task, or the price that the principal has to beat for the agent to complete the task for them. when setting a contract. To capture the agent’s additional information and opportunity to improve the metrics by which they are evaluated, our model supposes that the agent is privately aware of additional variables that correlate with their cost of task completion, and further has the opportunity to reveal these additional variables to the principal prior to the design of the contract. We analyze when the agent prefers a contract that depends on the revealed metrics over a contract that ignores them.
Importantly, our model deviates from the standard Bayesian setup in persuasion games, since without the agent’s help, the principal is not even aware of the existence of the metric. We also assume the agent cannot lie about or selectively mask the metric depending on its value, reflecting a common reality where online platforms and regulators often have significant data collection power, and can even compel an evaluated agent to report metrics [Dranove et al., 2003].
@@
While the principal is always better off knowing more metrics, the incentives for the agent are nuanced—revealing a metric reduces the amount of information rent the agent can extract when their costs turn out to be low. However, better information also means more high-cost tasks will go forward that the agent otherwise would not have accepted, as the optimal contract would adequately reward the agent when the metric indicates a high cost.
We prove that the agent prefers to reveal metrics that strongly differentiate the highest cost settings from rest, and conceal metrics that strongly differentiate the lowest cost settings from the rest. For example, a university might want to reveal to a government funder the number of low-income students that matriculate (indicating higher operating costs), but conceal the amount of funding they anticipate from private donations (indicating that they could continue without additional funds).
Our analysis becomes richer when we expand the agent’s action space to include the ability to garble or add noise to their metric before revealing it. We analyze a garbling mechanism that guarantees a notion of local differential privacy for the metric, and show that under a fairly wide set of conditions, the agent may prefer to reveal a garbled metric over both fully concealing and fully revealing the original metric. In fact, garbling can lead to Pareto improvement for both the principal and agent. From a policy standpoint, this demonstrates settings when both a principal and agent can derive economic value from a privacy preserving mechanism, even before considering the inherent social value of privacy.
Finally, we demonstrate an application of our theory to analyzing feature discovery for pricing on rideshare platforms, using public data scraped from Uber and Lyft.
@@
    We introduce a model for the elicitation of unknown metrics, via an agency game with information transfer.
@@
    We present sufficient conditions under which an agent would prefer to reveal or conceal a metric.
@@
    We show that when the agent has the ability to reveal a garbled metric, this can lead to Pareto improvement for both the principal and agent.
@@
    We leverage connections between our model and price discrimination to analyze total welfare.
  5. 5.
    We apply our model with real data to analyze feature discovery for pricing on rideshare platforms.
2 Related Work
Our model builds on literature from contract design and agency games. It also fits into a large literature on information design, and can be seen as an instance of information design with a specific structure. Our model also overlaps with a large body of work on price discrimination.
Agency games and contract design. We build from the well established contract design problem of Laffont and Tirole [1986], which concerns a principal’s design of a contract when an agent’s effort and cost type are privately held by the agent. Key to this setting are asymmetries in values and information between the principal and agent, and the literature explores issues of moral hazard and adverse selection that arise from these asymmetries [Laffont and Martimort, 2009, Gibbons et al., 2013, Milgrom and Roberts, 1992].
@@ 2 Related Work
Persuasion games and information design. The use of stakeholder-supplied information for decision-making was introduced by Milgrom and Roberts [1986] in their seminal work on persuasion games. Milgrom and Roberts [1986] give an example of analyzing a buyer’s purchasing strategy when a seller can send a quality signal about their product. While this broad motivation of information transfer from interested parties is close to our setting of eliciting metrics from evaluated agents, our model has a key distinction that without the agent’s help, the principal could not even access a prior over the metric. We also constrain the agent’s information revelation strategy to not be able to depend on the realized value of the variable. This brings our setting closer to a problem of metric discovery, where the fundamental problem is a principal’s lack of awareness of a metric’s observability, rather than a realized signal value [Spence, 1978].
@@
More broadly, Bayesian persuasion and information design provides a general framework for analyzing the effects of the distribution of information on the outcomes of a game [Kamenica, 2019, Bergemann and Morris, 2019]. Our model can be seen as a specific instantiation of an information design problem where the sender is the agent, the receiver is the principal, and the contracting relationship determines the principal’s and agent’s action spaces and equilibria. We impose a constraint over the sender’s information transfer policy in order to capture the incentives for an agent to reveal observability of a variable to the principal. Notably, the agent cannot selectively mask or lie about their signal based on its realized value, as our model is motivated by settings where a principal has powerful data collection capabilities.
@@
Price discrimination. Our model has analogies with classical price discrimination, and thereby opens new avenues for applying the well-developed tools from price discrimination to understanding the metric design problem. It also uncovers new challenges that extend existing price discrimination perspectives. These issues have been anticipated by Bergemann et al. [2015], who define a mapping from third-degree price discrimination onto the class of agency problems and establish the existence of market segmentations that achieve all possible trade-offs between consumer and producer surplus within some basic constraints. Our work complements this analysis—instead of analyzing all possible segmentations, we consider the agent’s information-revelation incentives as a function of the properties of a specific segmentation induced by some cost-correlated variable which is initially only available to the agent.
@@
Our garbling setting also notably differs from the models of price discrimination with transportation costs and arbitrage [Wright, 1993] or restricted price discrimination [Aguirre et al., 2010]. While restricted price discrimination mechanisms also effectively interpolate between full discrimination and no discrimination, we prove a substantive difference between garbling and these mechanisms in Appendix G.5.
@@
Sunspots and correlated equilibria. Our model is also connected to the so-called sunspots literature [Woodford, 1990, Howitt and McAfee, 1992]. In this literature, there are multiple steady-state equilibria and an observable random variable produces a correlated equilibrium, with agents conditioning their behavior on this otherwise extraneous variable not because it matters to payoffs, but because it predicts others’ behaviors. The literature is known as sunspots because, during the 19th century, some people believed sunspot activity predicted agricultural yields [Jevons, 1884], and while it did not, it could predict the behavior of commodity traders who believed it did. This literature is the polar opposite of the problem we study: sunspots are a known, extraneous variable believed to be relevant, while we study an unknown (to the principal) relevant variable.
@@
Preferences for privacy. Our model interacts with literatures on privacy and information-hiding, by highlighting situations where agents experience conflicting incentives both for and against sharing information. The result can be that agents may prefer partial sharing, which echoes the subtleties that arise elsewhere when privacy and behavior interact [Cummings et al., 2016, Acquisti et al., 2016].
Several works have considered the interaction between privacy and price discrimination in particular [Acquisti and Varian, 2005, Conitzer et al., 2012, Montes et al., 2015, Fudenberg and Villas-Boas, 2006, Ananthakrishnan et al., 2024b, Fallah et al., 2024]. Close to our setting is Fallah et al. [2024], who consider a noise addition mechanism that distorts the principal’s view of the observed value distributions per market segment. However, unlike our model, they assume that the principal does not know the aggregate market (the marginal distribution of costs in our setting), and their noise model can distort the principal’s belief about the aggregate market. Our garbling mechanism also differs from these other works in various ways (including, e.g., that there is no cost to noise addition, that the agent does not derive inherent value from garbling, and that noise is only added to the metric). Still, these various models have led to similar findings that privacy can yield economic benefits.
3 Agency Game with Information Transfer
@@
Specifically, suppose a principal contracts the agent to complete a task, where an agent may exert binary effort. Suppose the principal receives value $b$ if the agent exerts effort and completes the task, and 00 otherwise (we assume that the task is completed deterministically if the agent exerts effort). Suppose the agent incurs cost $C\in\mathbb{R}_{+}$ for exerting effort. The exact cost is unobserved to the principal, but the principal is aware of a prior distribution over agent’s cost type, denoted by the random variable $C$. The agent observes both the proposed price and their realized cost type before deciding whether or not to exert effort. This maps onto the well-established agency game setup where the principal must design a contract when the agent’s cost and effort are private [Laffont and Tirole, 1986].
To model the agent’s additional knowledge of a metric, suppose the agent is aware of a variable $X\in\mathcal{X}$ which is correlated with their cost $C$. However, $X$ is an unknown unknown to the principal. Prior to the principal’s design of the contract, the agent has a choice of whether or not to reveal $X$, which entails revealing both the observability of $X$ at the time of the design of the contract, and the realized value of $X$ at the execution of the contract (which the principal can verify). The fact that $X$ is not observable to the principal without the agent revealing is a key distinction between our model and prior work in signaling games. That is, the principal cannot rely on prior information over $C$ and $X$, since they are not even aware of the existence of $X$. This allows us to analyze a decision problem that differs from that of the standard Bayesian framework, which instead focuses on an agent’s decision to reveal their value of $X$ after it is realized.
If the agent chooses to not reveal $X$, then the rest of the game proceeds as a standard agency game with private cost: the principal designs a contract based on their knowledge of the prior distribution $C$. If the agent reveals $X$, then the principal can offer a contract that conditions on the realized value of $X$. The timing of the full agency game with the possibility of agent information transfer is summarized in Figure 1. The text in black matches the standard agency game [Laffont and Tirole, 1986], and the text in gray represents additional elements introduced by our model.
@@
The principal’s contract design problem when $X$ is concealed reduces to choosing a single price $p$ where the agent is paid $p$ if the task is completed, and zero otherwise. If $X$ is revealed, then the principal offers a contract with distinct prices $\rho(x)$ for different realized values of $x\in\mathcal{X}$. At the execution of the contract, the agent receives payment $\rho(x)$ if the task is completed and $X=x$, and 00 otherwise. We assume that the principal still receives the same value $b$ if the task is completed, regardless of $X$.
The key question in this work concerns the agent’s decision of whether or not to reveal $X$ at time $t=1$. In Section 4, we treat this as a binary decision of whether or not to reveal $X$; we will later relax this in Section 5 to expand the agent’s action space to reveal a garbled version of $X$, thus interpolating between the concealed and revealed settings.
3.1 Optimal Contracts
We begin by outlining the optimal contract and equilibrium utilities of the principal and agent when the contract is agnostic of $X$, which we refer to as the concealed information setting. Then, we outline the optimal contract and equilibrium utilities when the principal can condition on $X$ to determine payments to the agent, which we refer to as the revealed information setting. We assume that both principal and agent are risk neutral throughout.
Concealed information contract. If the metric $X$ is not revealed at $t=1$, then the rest of the game proceeds as a standard agency game with private cost, where the principal chooses a single transfer $p$ based on their knowledge of the prior distribution over the agent’s cost $C$. The agent’s optimal policy at the execution of the contract is to exert effort if their realized cost is less than or equal to the payment. Thus, the agent’s best response to the principal’s choice of transfer $p$ is to exert effort with probability $F(p)={\mathbb{P}}(C\leq p)$ (the set where $C=p$ has measure zero). [^2] We may also think of $F(p)$ as the task completion “quantity” as a function of price $p$, or the proportion of agents drawn uniformly at random from a population with costs distributed as $C$ that would complete the task for price $p$.
$\Pi_{\text{con}}(p)\coloneqq F(p)(b-p).$
The agent’s expected utility under the principal’s choice of transfer $p$ is given by
$V_{\text{con}}(p)\coloneqq\mathbb{E}[(p-C)\operatorname{\mathbbm{1}}(C<p)].$
The principal moves first and chooses
$p^{*}\in\arg\max_{p\geq 0}\Pi_{\text{con}}(p).$
The agent’s utility at equilibrium is then $V_{\text{con}}(p^{*})$.
Revealed information contract. If the metric $X$ is revealed at time $t=1$, then the principal can vary the payment amount depending on $X$, denoted as $\rho:\mathcal{X}\to\mathbb{R}_{+}$. The agent’s best response to the principal’s payment function $\rho(\cdot)$ is to exert effort with probability $F_{x}(\rho(x))=P(C\leq\rho(x)|X=x)$ when $X=x$. The principal’s expected utility in the revealed setting is given by
$\Pi_{\text{rev}}(\rho)\coloneqq\mathbb{E}[F_{X}(\rho(X))(b-\rho(X))].$
The agent’s expected utility under the principal’s choice of transfer function $\rho(\cdot)$ is
$V_{\text{rev}}(\rho)\coloneqq\mathbb{E}[(\rho(X)-C)\operatorname{\mathbbm{1}}(C<\rho(X))].$
The principal again moves first and chooses
$\rho^{*}\in\arg\max_{\rho\in\mathcal{F}}\Pi_{\text{rev}}(\rho).$
The agent’s choice of whether to reveal or conceal is ultimately determined by their utility difference, $V_{\text{rev}}(\rho^{*})-V_{\text{con}}(p^{*})$. The agent prefers to reveal $X$ if and only if the utility difference is positive.